On January 1, 2018, Solo Inc. issued 1,200 of its 7%, $1,000 bonds at 97.8. Interest is payable semiannually on January 1 and July 1. The bonds mature on January 1, 2028. Solo paid $52,000 in bond issue costs. Solo uses straight-line amortization.

What is the carrying value of the bonds reported in the December 31, 2018, balance sheet?
Answer #1

Solo Inc

  1. Determination of the carrying value of the bonds reported in the December 31, 2018 balance sheet:

Carrying value of the bonds reported in December 31, 2018 balance sheet = $1,181,440

Cash received on bond issue = 1,200 x 978 = $1,173,600

Less: bond issue costs = $52,000

Net cash received = $1,121,600

Discount on bond issue –

Face value of bond issue = 1,200 x $1,000 = $1,200,000

Add: bond issue costs = $52,000

Discount on bond issue = $78,400

Period of bond issue = 10 years (2028 – 2018)

Discount amortization = $78,400/10 = $7,840

Annual cash interest = 1,200,000 x 7% = $84,000

Interest expense = 84,000 + 7,840 = $91,840

Carrying value of bond as on January 1, 2018 = $1,173,600

Add: Discount on bond amortization = $7,840

Carrying value of bond as on December 31, 2018 = 1,181,440